OpenStudy (anonymous):

What are the domain and range of the function y=2√3x+4 -5

OpenStudy (anonymous):

you're having trouble with these arent you? :P its pretty easy once you get the hang of it. for this domain can you do this same thing i showed you the last question? set it up like so \[3x+4\ge0\]

OpenStudy (anonymous):

3x+4≥0 -4 -4 3x≥-4 3 3 x≥-1.33333

OpenStudy (anonymous):

Since we will use for finding the range it is simplier to keep it as \[\frac{ -4 }{ 3 }\]

OpenStudy (anonymous):

I'm just not good at setting up the first part. ok

OpenStudy (anonymous):

thanks.

OpenStudy (anonymous):

alright :) now simply plug in -4/3 for x in the equation to find the range

OpenStudy (anonymous):

-5

OpenStudy (anonymous):

yes! good job. that is the range :)

OpenStudy (anonymous):

OK. Thank you so much!

OpenStudy (anonymous):

Could you help me with this problem? When will the dependent variable in the equation y=√x+4 -3 equal or exceed 4?

OpenStudy (anonymous):

x is often referred to as the independent variable, and y is referred to as the dependent variable. This is because y is a function of x. Now, domain is all the x-values, and the range is all the y-values. So I assume it is asking for when the range will equal or exceed 4. One second, ill set it up for you.

OpenStudy (anonymous):

\[\sqrt{x+4}-3=4\] \[\sqrt{x+4}=7\] now just solve the equation :) i did the first two steps for you.

OpenStudy (anonymous):

whoops that should be \[\ge \]

OpenStudy (anonymous):

x ≥ 53 ?

OpenStudy (anonymous):

uh oh. you must've missed a step. lets go over it. \[\sqrt{x+4}=7\] Alright, how do you get rid of a radical?

OpenStudy (anonymous):

( )^2

OpenStudy (anonymous):

yep! square both sides.

OpenStudy (anonymous):

(x+4)^2= x+4 (7)^2=49

OpenStudy (anonymous):

remember you are aiming to get x by itself!

OpenStudy (anonymous):

x^2

OpenStudy (anonymous):

x+4=49

OpenStudy (anonymous):

-4 -4 x=45

OpenStudy (anonymous):

x ≥ 45

OpenStudy (anonymous):

yep! do you understand why? i did keep forgetting that greater than/equal to sign, my bad.

OpenStudy (anonymous):

Ya, I understand now. Thank you for getting me started and helping me through it!! That's alright, I knew what you were saying.

OpenStudy (anonymous):

yay :) my pleasure. I love these kinds of problems. good luck!